# Exploring Diffusion Transformer Designs via Grafting

### v1

@@ 1 Introduction
The space of architecture editing is vast, raising a practical question: what types of replacements should we study? To ground this investigation, we first establish a self-grafting baseline that replaces existing operators with randomly initialized weights. We show that our two-stage grafting procedure recovers near-baseline model quality, validating the approach. Building on this, we focus on replacing existing operators with efficient alternatives, aiming to reduce model FLOPs while preserving quality. In some cases, we explore replacements that increase model FLOPs to examine broader design choices. To study this systematically, we construct a testbed based on DiT-XL/2. We define a set of architectural edits that enable us to evaluate how different grafting schemes affect model quality. We highlight our design axes: (1) which operator to replace (e.g., MHA, MLP); (2) what to replace it with (e.g., convolutions); (3) how to select layers to edit (e.g., all layers); (4) replacement ratio (full vs. partial). The motivation for replacing MHA and MLP operators stems from both empirical evidence and architectural considerations: for MHA, our attention-locality analysis motivates the use of local operators, while for MLP, we adopt prior architecture ideas fu2023monarch ; komatsuzakisparse ; kaplan2020scaling .
@@ 4.1 Testbed and Experiment setup
We construct a testbed to systematically evaluate how design decisions affect generative quality under grafting. We focus on efficient replacements that reduce FLOPs, but also include higher-FLOP variants to explore a broader range of architectural edits. We target MHA and MLP operators, which account for a significant portion of FLOPs in DiTs compared to other operators (e.g., normalization, activation, residuals). The rationale for replacing MHA or MLP operators is grounded in both empirical and architectural considerations: for MHA, our attention locality analysis motivates the use of local operators; for MLP, we leverage prior architecture ideas fu2023monarch ; kim2024solar ; qiu2024compute ; petty2024impact ; komatsuzakisparse ; kaplan2020scaling . Given a replacement operator, the decision to graft it to a model with $L$ transformer layers spans a space of $2^{L}$ configurations. To make this tractable, we study two layer selection strategies: full (replace all operators) and interleaved (replace operators in a repeating pattern) strategies. The latter is inspired by striped transformer designs poli2024mechanistic ; ku2025systems ; Brixi2025.02.18.638918 . Our testbed is detailed in Tab. 3.
@@
Experiment setup. For our hybrid experiments, we mostly use the hyperparameters determined from our self-grafting studies.
@@ 4.2 Results and Insights
Table 4: Generation quality and efficiency of grafted hybrid architectures in DiT-XL/2. We report quality (IS, FID, sFID, Precision, Recall) and efficiency ($\Delta$FLOPs and $\Delta$Param) results for MHA and MLP grafting experiments. For each alternative, setups that maintain FID within 0.5 of the baseline and offer the largest FLOPs reduction (or smallest FLOPs increase) are highlighted.  denotes setups with poor generation (FID > 50). $\Delta\text{FLOPs}$ and $\Delta\text{Param}$ denote the percentage change in FLOPs and parameters, respectively (for MHA and MLP operators). For MHA, total cost is split into $\Delta\text{FLOPs}_{\text{op.}}$ (softmax attention, gating) and $\Delta\text{FLOPs}_{\text{ft.}}$ (QKV/output projections, featurizers). We do not use this decomposition for MLP variants. Mamba-2 incurs higher $\Delta\text{FLOPs}_{\text{ft.}}$ due to additional projections. Exact FLOP calculations are provided in Sec. F.1. Key result: Many interleaved designs achieve good quality generation (FID within 0.5 of baseline).
@@ 6 Case Study: Converting Model Depth to Width via Grafting
Experiment Setup. The rewiring schematic is shown in Fig. 5. Stage 1: Activation distillation. Each parallel pair was initialized via activation distillation using L1 regression. The weights for each block in the parallel pair were initialized from their corresponding pre-trained weights, rather than random initialization. Similar to our previous experiments, 8k ImageNet-1K samples were used for this stage. Stage 2: Lightweight finetuning. Given the architectural restructuring, finetuning was performed using 25% of the training data. The learning rate was linearly warmed up to 1e-4 and halved at 75k and 150k iterations. Experiment details can be found in Sec. C.
@@
Takeaway 3: Grafting enables architectural restructuring at the transformer block level, allowing depth to be traded for width.
@@
Table 6: Generative quality vs. model depth. Generative quality metrics are reported: FID, IS, sFID, Precision, and Recall. A.R. (Aspect Ratio) is defined as model width divided by depth (e.g., 1152/14 = 82.3). Parameters (Params) are reported in millions. For pruning and grafting setups, we report speedup with respect to DiT-XL/2 (depth=28). Off-the-shelf DiT-L/2, U-ViT-L, and DiT-B/2 scores, along with pruning baselines (BK-SDM, TinyDiT-D14, and TinyDiT-D14 w/ MKD), are sourced from  fang2024tinyfusion . MKD refers to Masked Knowledge Distillation, a recovery method used in TinyDiT fang2024tinyfusion . Our grafted models achieve better generative quality at depth=14, surpassing baselines in FID, IS, Precision, and Recall. Speedup is measured for a single forward pass on an Nvidia H100 (batch size=2). More implementation details are provided in Sec. C.
@@ 7 Related Work
Architectural editing of pretrained models. Another line of work focuses on linearizing large language models by replacing softmax attention with efficient operators, such as linear attention zhanglolcats ; wang2024the ; bick2024transformers . Similar ideas have also been adopted for diffusion models in  liu2024clear ; liu2024linfusion ; becker2025edit , though these works focus only on ultra-high-resolution settings. These prior efforts typically focus on replacing a single operator type (primarily attention) or are limited to specific application domains. Grafting presents a more general and comprehensive approach for architectural editing. It extends beyond single-operator replacement to enable modifying multiple operator types, exploring diverse architectural alternatives (e.g., both MHA and MLP replacements), and restructuring entire architectures (e.g., converting depth to width). Recently, FFN Fusion bercovich2025ffn explored parallelizing transformer blocks in LLMs, aiming to reduce sequential computation.
@@ 8 Conclusion and Discussion
Limitations. This work primarily focuses on architectural editing of pretrained Diffusion Transformers (DiTs), specifically targeting Multi-Head Attention (MHA) and Multi-Layer Perceptron (MLP) components. Other architectural components, such as normalization layers and activation functions, will be explored in future work. We note that these are latent diffusion models, and grafting components in their corresponding VAEs remains an area for future study. Our experiments primarily focus on DiTs, and generalizing grafting to other model families, such as autoregressive models, is a direction for future research. The PixArt-$\Sigma$ setup used synthetic data for grafting, which may propagate artifacts and biases into the grafted models. While this work focuses on architectural editing, it remains an open question whether architectures that perform well under grafting also perform well when trained from scratch. Finally, grafting requires access to a pretrained model.
Applications and future work. Grafting holds promise for diverse applications where efficiency is important. This includes adapting models from low-resolution to high-resolution generation, extending capabilities from short-form video understanding/generation to long-form chandrasegaran2024hourvideo ; chen2025eagle , or improving user experience in interactive applications like image editing where even modest speedups (e.g., 10%) are highly valued. We hope that our testbed, insights, and results will encourage the community to actively explore new architecture designs.

### v2

@@ 1 Introduction
The space of architectural editing is vast, raising a practical question: what types of replacements should we study? We first establish a self-grafting baseline, where we replace all MHA and MLP operators in DiT-XL/2 with randomly initialized counterparts. Despite the scale of this intervention, our grafting procedure achieves near-baseline model quality using under 1% of pretraining compute. Building on this, we focus on replacing existing operators with efficient alternatives, aiming to reduce model FLOPs while preserving quality. We also explore replacements that increase model FLOPs to examine broader design choices. To study this systematically, we construct a testbed based on DiT-XL/2 and define a set of architectural edits to evaluate how different grafting schemes affect model quality. We organize our design space along four axes: (1) which operator to replace (e.g., MHA, MLP); (2) what to replace it with (e.g., convolutions); (3) how to select layers for replacement (e.g., all layers); and (4) replacement ratio (full vs. partial). We focus on replacing MHA and MLP operators, as they account for a large fraction of model FLOPs. Replacements for MHA and MLP operators are motivated by empirical findings and prior architectural designs: our locality analysis supports local operators for MHA, while for MLP, we adopt ideas from prior work fu2023monarch ; komatsuzakisparse ; kaplan2020scaling .
@@ 4.1 Testbed and Experiment setup
We construct a testbed to systematically evaluate how design decisions affect generative quality under grafting. We focus on efficient replacements that reduce FLOPs, but also include higher-FLOP variants to explore a broader range of architectural edits. We target MHA and MLP operators, which account for a significant portion of FLOPs in DiTs compared to other operators (e.g., normalization, activation, residuals). The rationale for replacing MHA or MLP operators is grounded in both empirical and architectural considerations: for MHA, our attention locality analysis (Fig. 2) motivates the use of local operators; for MLP, we leverage prior architecture ideas fu2023monarch ; kim2024solar ; qiu2024compute ; petty2024impact ; komatsuzakisparse ; kaplan2020scaling . Given a replacement operator, the decision to graft it to a model with $L$ transformer layers spans a space of $2^{L}$ configurations. To make this tractable, we study two layer selection strategies: full (replace all operators) and interleaved (replace operators in a repeating pattern) strategies. The latter is inspired by striped transformer designs poli2024mechanistic ; ku2025systems ; Brixi2025.02.18.638918 . Our testbed is detailed in Tab. 3.
@@
Experiment setup. For our hybrid experiments, we mostly use the hyperparameters determined from our self-grafting studies (Sec. 3.2).
@@ 4.2 Results and Insights
Table 4: Generation quality and efficiency metrics for MHA and MLP grafting. We report quality (IS, FID, sFID, Precision, Recall) and efficiency ($\Delta$FLOPs and $\Delta$Param) results. Baseline refers to DiT-XL/2. For each alternative, setups that maintain FID within 0.5 of the baseline and offer the largest FLOPs reduction (or smallest FLOPs increase) are highlighted.  denotes setups with poor generation (FID > 50). $\Delta\text{FLOPs}$ and $\Delta\text{Param}$ denote the percentage change in operator FLOPs and parameters, respectively. For MHA, total cost is split into $\Delta\text{FLOPs}_{\text{op.}}$ (softmax attention, gating) and $\Delta\text{FLOPs}_{\text{ft.}}$ (QKV/output projections, featurizers). We do not use this decomposition for MLP variants. Mamba-2 incurs higher $\Delta\text{FLOPs}_{\text{ft.}}$ due to additional projections. FLOP expressions are provided in Sec. F.1. Key result: Many interleaved designs achieve good quality generation (FID within 0.5 of baseline). All experiments use 10% training data and $<$1% pretraining compute.
@@ 6 Case Study: Converting Model Depth to Width via Grafting
Experiment Setup. The rewiring schematic is shown in Fig. 5. We use DiT-XL/2. Stage 1: Activation distillation. Each parallel pair was initialized via activation distillation using L1 regression. The weights for each block in the parallel pair were initialized from their corresponding pre-trained weights, rather than random initialization. Similar to our previous experiments, 8k ImageNet-1K samples were used for this stage. Stage 2: Lightweight finetuning. Given the architectural restructuring, finetuning was performed using 25% of the training data. The learning rate was linearly warmed up to 1e-4 and halved at 75k and 150k iterations. Additional details can be found in Sec. C.
@@
Takeaway 3: Grafting enables architectural restructuring at the transformer block level, allowing model depth to be traded for width.
@@
Table 6: Generative quality vs. model depth. We report generative quality metrics (IS, FID, sFID, Precision, and Recall). A.R. (Aspect Ratio) is defined as model width divided by depth (e.g., 1152/14 = 82.3). Parameters (Params) are reported in millions. For pruning and grafting setups, we report speedup with respect to DiT-XL/2 (depth=28). Off-the-shelf DiT-L/2, U-ViT-L, and DiT-B/2 scores, along with pruning baselines (BK-SDM, TinyDiT-D14, and TinyDiT-D14 w/ MKD), are sourced from  fang2024tinyfusion . MKD refers to Masked Knowledge Distillation, a recovery method used in TinyDiT fang2024tinyfusion . Speedup is measured for a single forward pass on an Nvidia H100 (batch size=2). More implementation details are provided in Sec. C. Key Result. Our grafted models achieve better generative quality at depth=14, surpassing baselines in FID, IS, Precision, and Recall.
@@ 7 Related Work
Architectural editing of pretrained models. Another line of work focuses on linearizing large language models by replacing softmax attention with efficient operators, such as linear attention zhanglolcats ; wang2024the ; bick2024transformers . Similar ideas have also been adopted for diffusion models in  liu2024clear ; liu2024linfusion ; becker2025edit , though these works focus only on ultra-high-resolution settings. These prior efforts typically focus on replacing a single operator type (primarily attention) or are limited to specific application domains. Grafting presents a more general and comprehensive approach for architectural editing. It extends beyond single-operator replacement to enable modifying multiple operator types, exploring diverse architectural alternatives (e.g., both MHA and MLP replacements), and restructuring architectures (e.g., converting model depth to width). Recently, FFN Fusion bercovich2025ffn explored parallelizing transformer blocks in LLMs, aiming to reduce sequential computation.
@@ 8 Conclusion and Discussion
Limitations. This work primarily focuses on architectural editing of pretrained Diffusion Transformers (DiTs), specifically targeting self-attention and MLP operators. Other architectural components, such as normalization layers and activation functions, will be explored in future work. We note that these are latent diffusion models, and grafting components in their corresponding VAEs remains an area for future study. Our experiments primarily focus on DiTs, and generalizing grafting to other model families, such as autoregressive models, is a direction for future research. The PixArt-$\Sigma$ setup used synthetic data for grafting, which may propagate artifacts and biases into the grafted models. While this work focuses on architectural editing, it remains an open question whether architectures that perform well under grafting also perform well when trained from scratch. Finally, grafting requires access to a pretrained model.
Applications and future work. Grafting holds promise for diverse applications where efficiency is important. This includes adapting models from low-resolution to high-resolution settings, extending capabilities from short-form video understanding/generation to long-form chandrasegaran2024hourvideo ; chen2025eagle , or improving user experience in interactive applications like image editing where even modest gains (e.g., 10% speedup) are highly valued. We hope that our testbed, insights, and results will encourage the community to actively explore new architecture designs. Code and grafted models: grafting.stanford.edu.